valentinashehu3340:

A customer from Cavallaro's Fruit Stand picks a sample of 5 oranges at random from a crate containing 65 oranges, of which 6 are rotten. What is the probability that the sample contains 1 or more rotten oranges? (Round your answer to three decimal places

valentinashehu3340:

Help

mikewwe13:

To solve this problem, we can use the complement rule. The probability that the sample contains 1 or more rotten oranges is equal to 1 minus the probability that the sample contains no rotten oranges. The probability of picking one orange that is not rotten is (65 - 6) / 65 = 59 / 65. The probability of picking five oranges that are not rotten is (59/65)^5. Therefore, the probability of the sample containing no rotten oranges is (59/65)^5. Hence, the probability of the sample containing 1 or more rotten oranges is: 1 - (59/65)^5 = 0.461 (rounded to three decimal places) Therefore, the probability that the sample contains 1 or more rotten oranges is 0.461.